What is the electric potential at the centre of square of side 1m? The charges 1×10^-8 C , -2×10^-8 C, 3×10^-8C and 2×10^-8 C are placed at the corners of the square.

Grade:11

1 Answers

Khimraj
5 years ago
potential at center of square is equal to alzebric sum of potential at center due to individual charge.
V = K(1-2+3+2)*10-8/(1/\sqrt{2}) = 360\sqrt{2} V
Hope it clears.........................................